Two persons A and B appeared in an interview for two vacancies for the same post. The probability of A’s selection is 1/7 and that of B is 1/5. What is the probability that
i. Both of them are selected
ii. Only one of them is selected

Answer by stanbon(75887)

Two persons A and B appeared in an interview for two vacancies for the same post. The probability of A’s selection is 1/7 and that of B is 1/5. What is the probability that
i. Both of them are selected
P(A and B) = (1/5)(1/7) = 1/35
------------------------------------
ii. Only one of them is selected
P(one selected) = 1 - [P(both selected) + P(neither is selected)]
= 1 - [1/35 + (4/5)(6/7)]
= 1 - [1/35 + 24/35]
= 1 - (5/7)
= 2/7
=================
Cheers,
Stan H.